OpenStudy (anonymous):

four times a number added to 7 times the number = 33 , find the number.

OpenStudy (anonymous):

4x+7=33 where x is the unknown no. now solve fr x

OpenStudy (anonymous):

x=1.0 ?

OpenStudy (anonymous):

or 3 ? im not sure im lost

OpenStudy (anonymous):

its 4x+7x=33

OpenStudy (anonymous):

srry

OpenStudy (anonymous):

so 11x=33 x=3

OpenStudy (anonymous):

can you help me with another one?

OpenStudy (anonymous):

anything for pretty girls

OpenStudy (anonymous):

when 4 times a number is subtracted from 7 times the number , the result is 27, find the number

OpenStudy (anonymous):

im coming up with 9 ?

OpenStudy (anonymous):

7x-4x=27 SOLVE

OpenStudy (anonymous):

3x=27 x=9

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

anytime

OpenStudy (anonymous):

u a freshman?

OpenStudy (anonymous):

sophmore college, have to take this dumb class! hate math

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

dont wrry